# RichControl: Structure- and Appearance-Rich Training-Free Spatial Control for Text-to-Image Generation

### v3

@@ 5.2 Comparison with State-of-the-Art (SOTA)
Figure 9: Integrating our method as a plug-in into FreeControl [52]. Our method consistently improves FreeControl on both quantitative and qualitative evaluations, achieving stronger structure fidelity and perceptual quality.
Figs. 5 and 1 present quantitative and qualitative comparisons between our method and existing baselines, respectively. While training-based approaches like ControlNet [102] and T2I-Adapter [53] exhibit lower Self-sim scores, they often fail to adhere to the text prompts (e.g., origami, white egg, sculpture in Fig. 5), leading to impaired text-image alignment. In contrast, our method achieves robust structural alignment while maintaining superior text-image consistency.
On the other hand, training-free baselines exhibit several limitations. SDEdit [51], PnP [80], and P2P [28] are prone to condition leakage, producing outputs that closely resemble the condition image. FreeControl [52] and InfEdit [97] yield unstable results, often generating images with inferior text-image alignment and artifacts (rows 1-4). Ctrl-X [47] performs reliably in many cases but still suffers structural misalignment (row 4), condition leakage (row 2), and artifacts (rows 1 and 3) as shown in Fig. 5. In contrast, our method consistently outperforms these baselines in structural preservation, text-image alignment, and visual quality, excelling in difficult scenarios such as abstract conditions (e.g., pose), multi-object scenes, and challenging prompts (e.g., origami). Quantitative evaluations further confirm our advantage. As shown in Fig. 1, our method surpasses training-free baselines across nearly all metrics. Please refer to Figs. 16 and 17 and Tab. 2 in the appendix for additional results.
Similar to prior studies [22, 95, 89, 50], we find that reward models exhibit stronger alignment with human preferences than traditional metrics. Accordingly, in the subsequent experiments, we primarily report results based on three reward-model metrics.
@@ 5.4 Improvement over Prior Methods
Our pipeline can serve as a plug-in to enhance other conditional T2I approaches. As an example, we applied our method to FreeControl [52], another recent conditional T2I model. As shown in Fig. 9, our approach consistently improves FreeControl across three evaluation metrics. Qualitative comparisons in Fig. 9 further highlight that FreeControl with our method yields outputs with both stronger structural preservation and higher perceptual quality.
@@ Appendix B Preliminaries
The backward process generates images by iteratively denoising an initial Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(0,\mathbb{I})$, also known as diffusion sampling [30]. This process uses a parameterized denoising network $\mathbf{\epsilon}_{\theta}$ conditioned on a text prompt $\mathcal{P}$, where at time step $t$ we obtain a cleaner $\mathbf{x}_{t-1}$ as
@@
The iterative inference of diffusion enables people to guide the sampling process on auxiliary information. Guidance modifies Eq. 4 to compose additional score functions that point toward richer and specifically conditioned distributions [6, 18], expressed as
@@
where $\mathbf{g}$ is an energy function and $s$ is the guidance strength. In practice, $\mathbf{g}$ can range from classifier-free guidance (where $\mathbf{g}=\mathbf{\epsilon}$ and $y=\emptyset$, i.e. the empty prompt) to improve image quality and prompt adherence for T2I diffusion [29, 64], to arbitrary gradients computed from auxiliary models or diffusion features common to guidance-based controllable generation [6, 18, 52]. Thus, guidance provides the customizability on the type and variety of conditioning for controllable generation, as it merely requires a differentiable loss with respect to $\mathbf{x}_{t}$. However, the need for backpropagation during inference often leads to increased memory consumption and slower inference speed. Moreover, guidance-based methods often fail to capture fine structural details in controllable generation tasks.
@@
Many pretrained T2I diffusion models are text-conditioned U-Nets, which contain an encoder and a decoder that downsample and then upsample the input $\mathbf{x}_{t}$ to predict $\mathbf{\epsilon}$, with long skip connections between matching encoder and decoder resolutions [30, 64, 61]. Each encoder/decoder block contains convolution layers, self-attention layers, and cross-attention layers: The first two control both structure and appearance, and the last injects textual information. Thus, many training-free controllable generation methods utilize these layers, through direct manipulation [28, 80, 38, 2, 97] or for computing guidance losses [18, 52], with self-attention most commonly used. Let $\mathbf{f}_{l,t}\in\mathbb{R}^{HW\times c}$ be the diffusion feature with height $H$, width $W$, and channel size $c$ at time step $t$ right before attention layer $l$. Then, the self-attention operation is
@@ C.1 KL Divergence
To analyze the domain gap between natural images and condition images, we collect 20 natural images from the ImageNet-T2IR dataset from [80]. Then we use the ControlNet processor [102] to convert these natural images into 5 conditions (canny edge, depth map, normal map, HED edge, and scribble drawing), resulting in 100 natural-condition image pairs.
@@ C.2 Self-Similarity
Following [79], we adopt the DINO self-similarity distance [10] to quantify structural similarity between images. In Vision Transformer (ViT) [17], an image is first divided into a sequence of non-overlapping patches, which are then linearly embedded and processed as tokens. In each Transformer layer, the tokens are projected into queries, keys, and values as follows:
@@
As shown in [79], this self-similarity-based descriptor can effectively capture the structure of an image while ignoring appearance details. Given two images $\mathbf{I}_{1}$ and $\mathbf{I}_{2}$, their structural distance is computed as the $\ell_{2}$ distance between their self-similarity matrices:
@@ D.1 Spatially-Aware Appearance Transfer
We build our method on top of the spatially-aware appearance transfer mechanism proposed in Ctrl-X [47]. Specifically, given diffusion features $\mathbf{f}_{l,t}^{\text{out}}$ and $\mathbf{f}_{l,t}^{\text{app}}$ from the output and appearance branches respectively, Ctrl-X [47] computes a cross-image attention map as follows:
@@ D.2 Appearance-Rich Prompting
Directly using the original prompt for appearance transfer may lead to artifacts in the generated image, since such prompts tend to be brief and lacking in semantic correspondence with the condition image (see Sec. 4.3 for details). To overcome this limitation, we propose a pipeline that enriches the original text prompt $\mathcal{P}$ with semantic information extracted from the structure condition image $\mathbf{I}^{\text{struct}}$, yielding a more appearance-rich prompt $\mathcal{P}^{\text{app}}$ for generating the final appearance image $\mathbf{I}^{\text{app}}$. As illustrated in Fig. 12, we first utilize GPT-4o [1] to extract key semantic entities from the condition image to produce dictionary $\langle d1\rangle$. To facilitate semantic alignment between the condition image and the text prompt, we further employ GPT-4o [1] to identify and associate these extracted entities with semantically-related elements in the original text, modifying $\langle d1\rangle$ to produce $\langle d2\rangle$. Finally, we revise the original prompt $\mathcal{P}$ using the extracted semantic information, producing an enhanced appearance prompt $\mathcal{P}^{\text{app}}$. To help the multimodal LLM correctly follow instructions and mitigate erroneous semantic transfer, our pipeline stores intermediate information in structured dictionaries, enabling more controlled and interpretable prompt editing. More examples of the Appearance-Rich Prompting (ARP) module are provided in Fig. 14. For the full prompt used with GPT-4o [1] for Appearance-Rich Prompting, see the accompanying .txt file in the supplementary material.
@@ E.1 Implementation Details
We implement our method with Diffusers [81] on SDXL 1.0 [61] and adopt the same injection layers following previous work [47]. We sample $\mathbf{I}$ with 50 steps of DDIM sampling and set $\eta=1$ [72]. For structure-rich injection, we set $\tau=400$ and $C=600$. For restart refinement, we set $\sigma_{t_{\min}}=1.0$, $\sigma_{t_{\max}}=2.0$, $N=3$, $S=5$, where $S$ is the total number of timesteps in the restart backward process. For self-recurrence, we set $t_{\text{min}}^{\prime}=500$, $t_{\text{max}}^{\prime}=900$, $N^{\prime}=2$, where $t_{\text{max}}^{\prime}$ is the self-recurrence starting point, $t_{\text{min}}^{\prime}$ is the self-recurrence end point, and $N^{\prime}$ is the number of self-recurrence [47]. We run all experiments on NVIDIA Tesla V100 GPUs.
@@ E.2 Dataset Details
We construct our dataset based on the conditional generation datasets from Ctrl-X [47] and FreeControl [52]. Specifically, for conditions canny edge, depth map, normal map, HED edge, and scribble drawing, we select condition-prompt pairs from both datasets and merge them. We collect a total of 15 condition images per condition and form 22 condition-prompt pairs for canny edge and 21 pairs for each of the remaining four conditions.
For human pose and segmentation map, since both original datasets contain limited examples, we supplement them by collecting additional human pose images from the web and segmentation masks from the ADE20K [110] dataset. We obtain 15 images for each of these two conditions and pair them with text prompts using a combination of templates and hand annotation, resulting in 21 image-prompt pairs for human pose and 23 for segmentation mask.
@@ Appendix F Additional Experimental Results
ControlNet [102] 0.067 0.309 0.701 0.509 0.298 0.285
T2I-Adapter [53] 0.116 0.287 0.728 0.636 -0.050 0.261
SDEdit [51] 0.154 0.259 0.315 0.734 -1.374 0.189
P2P [28] 0.197 0.251 0.266 0.724 -1.786 0.168
PnP [80] 0.157 0.256 0.151 0.724 -1.789 0.168
InfEdit [97] 0.135 0.296 0.357 0.636 -0.202 0.244
FreeControl [52] 0.116 0.320 0.667 0.626 0.554 0.285
Ctrl-X [47] 0.104 0.315 0.650 0.579 0.291 0.283
@@ F.1 Inference Efficiency
Our framework inherits the computational efficiency of injection-based pipelines. Moreover, since $g(t)=C$ in our implementation, our pipeline is able to cache and reuse $\mathbf{f}_{l,g(t)}^{\text{struct}}$, in contrast to previous synchronous pipelines that perform inference at every timestep. Tab. 3 shows the average inference time of different settings in our experiments, with the refiner of SDXL [61] enabled. As shown in the table, structure-rich injection saves inference time by up to 9% with no noticeable degradation in generation quality. Our framework still achieves superior inference efficiency even with the restart refinement schedule, demonstrating simultaneous improvements in both generation quality and inference speed. Appearance-rich prompting module improves the robustness of the framework at the expense of an 8% inference time overhead. Overall, this training-free framework provides users with the flexibility to adapt to task-specific priorities and negligible overhead in most cases.
@@ F.3 Additional Quantitative Results
Since T2I-Adapter-SDXL [53] supports only four (canny, depth, normal, and pose) out of the seven condition types in our full dataset, we further conduct a quantitative comparison limited to these four types. As shown in Tab. 2, our method outperforms all baselines across almost every metric. Notably, these metrics jointly assess both structure preservation (e.g., DINO self-similarity [79], DreamSim [22]) and generation quality (e.g., ImageReward [95], HPSv2 [89]), highlighting the effectiveness of our approach.

### v4

@@ 5.2 Comparison with State-of-the-Art (SOTA)
Figure 9: Integrating our method as a plug-in into FreeControl [mo2023freecontrol]. Our method consistently improves FreeControl on both quantitative and qualitative evaluations, achieving stronger structure fidelity and perceptual quality.
Figs. 5 and 1 present quantitative and qualitative comparisons between our method and existing baselines, respectively. While training-based approaches like ControlNet [zhang2023controlNet] and T2I-Adapter [mou2023t2i] exhibit lower Self-sim scores, they often fail to adhere to the text prompts (e.g., origami, white egg, sculpture in Fig. 5), leading to impaired text-image alignment. In contrast, our method achieves robust structural alignment while maintaining superior text-image consistency.
On the other hand, training-free baselines exhibit several limitations. SDEdit [meng2021sdedit], PnP [tumanyan2023plug-and-play], and P2P [hertz2022prompt-to-prompt] are prone to condition leakage, producing outputs that closely resemble the condition image. FreeControl [mo2023freecontrol] and InfEdit [xu2023infedit] yield unstable results, often generating images with inferior text-image alignment and artifacts (rows 1-4). Ctrl-X [lin2024ctrlx] performs reliably in many cases but still suffers structural misalignment (row 4), condition leakage (row 2), and artifacts (rows 1 and 3) as shown in Fig. 5. In contrast, our method consistently outperforms these baselines in structural preservation, text-image alignment, and visual quality, excelling in difficult scenarios such as abstract conditions (e.g., pose), multi-object scenes, and challenging prompts (e.g., origami). Quantitative evaluations further confirm our advantage. As shown in Fig. 1, our method surpasses training-free baselines across nearly all metrics. Please refer to Figs. 16 and 17 and Tab. 2 in the appendix for additional results.
Similar to prior studies [fu2023dreamsim, xu2023imagereward, wu2023hpsv2, ma2025hpsv3], we find that reward models exhibit stronger alignment with human preferences than traditional metrics. Accordingly, in the subsequent experiments, we primarily report results based on three reward-model metrics.
@@ 5.4 Improvement over Prior Methods
Our pipeline can serve as a plug-in to enhance other conditional T2I approaches. As an example, we applied our method to FreeControl [mo2023freecontrol], another recent conditional T2I model. As shown in Fig. 9, our approach consistently improves FreeControl across three evaluation metrics. Qualitative comparisons in Fig. 9 further highlight that FreeControl with our method yields outputs with both stronger structural preservation and higher perceptual quality.
@@ Appendix B Preliminaries
The backward process generates images by iteratively denoising an initial Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(0,\mathbb{I})$, also known as diffusion sampling [ho2020ddpm]. This process uses a parameterized denoising network $\mathbf{\epsilon}_{\theta}$ conditioned on a text prompt $\mathcal{P}$, where at time step $t$ we obtain a cleaner $\mathbf{x}_{t-1}$ as
@@
The iterative inference of diffusion enables people to guide the sampling process on auxiliary information. Guidance modifies Eq. 4 to compose additional score functions that point toward richer and specifically conditioned distributions [bansal2023universalguidance, epstein2023selfguidance], expressed as
@@
where $\mathbf{g}$ is an energy function and $s$ is the guidance strength. In practice, $\mathbf{g}$ can range from classifier-free guidance (where $\mathbf{g}=\mathbf{\epsilon}$ and $y=\emptyset$, i.e. the empty prompt) to improve image quality and prompt adherence for T2I diffusion [ho2022cfg, rombach2022ldm], to arbitrary gradients computed from auxiliary models or diffusion features common to guidance-based controllable generation [bansal2023universalguidance, epstein2023selfguidance, mo2023freecontrol]. Thus, guidance provides the customizability on the type and variety of conditioning for controllable generation, as it merely requires a differentiable loss with respect to $\mathbf{x}_{t}$. However, the need for backpropagation during inference often leads to increased memory consumption and slower inference speed. Moreover, guidance-based methods often fail to capture fine structural details in controllable generation tasks.
@@
Many pretrained T2I diffusion models are text-conditioned U-Nets, which contain an encoder and a decoder that downsample and then upsample the input $\mathbf{x}_{t}$ to predict $\mathbf{\epsilon}$, with long skip connections between matching encoder and decoder resolutions [ho2020ddpm, rombach2022ldm, podell2023sdxl]. Each encoder/decoder block contains convolution layers, self-attention layers, and cross-attention layers: The first two control both structure and appearance, and the last injects textual information. Thus, many training-free controllable generation methods utilize these layers, through direct manipulation [hertz2022prompt-to-prompt, tumanyan2023plug-and-play, kim2023denseattn, alaluf2023crossimage, xu2023infedit] or for computing guidance losses [epstein2023selfguidance, mo2023freecontrol], with self-attention most commonly used. Let $\mathbf{f}_{l,t}\in\mathbb{R}^{HW\times c}$ be the diffusion feature with height $H$, width $W$, and channel size $c$ at time step $t$ right before attention layer $l$. Then, the self-attention operation is
@@ C.1 KL Divergence
To analyze the domain gap between natural images and condition images, we collect 20 natural images from the ImageNet-T2IR dataset from [tumanyan2023plug-and-play]. Then we use the ControlNet processor [zhang2023controlNet] to convert these natural images into 5 conditions (canny edge, depth map, normal map, HED edge, and scribble drawing), resulting in 100 natural-condition image pairs.
@@ C.2 Self-Similarity
Following [tumanyan2022splicing], we adopt the DINO self-similarity distance [caron2021dino] to quantify structural similarity between images. In Vision Transformer (ViT) [ViT], an image is first divided into a sequence of non-overlapping patches, which are then linearly embedded and processed as tokens. In each Transformer layer, the tokens are projected into queries, keys, and values as follows:
@@
As shown in [tumanyan2022splicing], this self-similarity-based descriptor can effectively capture the structure of an image while ignoring appearance details. Given two images $\mathbf{I}_{1}$ and $\mathbf{I}_{2}$, their structural distance is computed as the $\ell_{2}$ distance between their self-similarity matrices:
@@ D.1 Spatially-Aware Appearance Transfer
We build our method on top of the spatially-aware appearance transfer mechanism proposed in Ctrl-X [lin2024ctrlx]. Specifically, given diffusion features $\mathbf{f}_{l,t}^{\text{out}}$ and $\mathbf{f}_{l,t}^{\text{app}}$ from the output and appearance branches respectively, Ctrl-X [lin2024ctrlx] computes a cross-image attention map as follows:
@@ D.2 Appearance-Rich Prompting
Directly using the original prompt for appearance transfer may lead to artifacts in the generated image, since such prompts tend to be brief and lacking in semantic correspondence with the condition image (see Sec. 4.3 for details). To overcome this limitation, we propose a pipeline that enriches the original text prompt $\mathcal{P}$ with semantic information extracted from the structure condition image $\mathbf{I}^{\text{struct}}$, yielding a more appearance-rich prompt $\mathcal{P}^{\text{app}}$ for generating the final appearance image $\mathbf{I}^{\text{app}}$. As illustrated in Fig. 12, we first utilize GPT-4o [achiam2023gpt4] to extract key semantic entities from the condition image to produce dictionary $\langle d1\rangle$. To facilitate semantic alignment between the condition image and the text prompt, we further employ GPT-4o [achiam2023gpt4] to identify and associate these extracted entities with semantically-related elements in the original text, modifying $\langle d1\rangle$ to produce $\langle d2\rangle$. Finally, we revise the original prompt $\mathcal{P}$ using the extracted semantic information, producing an enhanced appearance prompt $\mathcal{P}^{\text{app}}$. To help the multimodal LLM correctly follow instructions and mitigate erroneous semantic transfer, our pipeline stores intermediate information in structured dictionaries, enabling more controlled and interpretable prompt editing. More examples of the Appearance-Rich Prompting (ARP) module are provided in Fig. 14. For the full prompt used with GPT-4o [achiam2023gpt4] for Appearance-Rich Prompting, see the accompanying .txt file in the supplementary material.
@@ E.1 Implementation Details
We implement our method with Diffusers [vonplaten2022diffusers] on SDXL 1.0 [podell2023sdxl] and adopt the same injection layers following previous work [lin2024ctrlx]. We sample $\mathbf{I}$ with 50 steps of DDIM sampling and set $\eta=1$ [song2020ddim]. For structure-rich injection, we set $\tau=400$ and $C=600$. For restart refinement, we set $\sigma_{t_{\min}}=1.0$, $\sigma_{t_{\max}}=2.0$, $N=3$, $S=5$, where $S$ is the total number of timesteps in the restart backward process. For self-recurrence, we set $t_{\text{min}}^{\prime}=500$, $t_{\text{max}}^{\prime}=900$, $N^{\prime}=2$, where $t_{\text{max}}^{\prime}$ is the self-recurrence starting point, $t_{\text{min}}^{\prime}$ is the self-recurrence end point, and $N^{\prime}$ is the number of self-recurrence [lin2024ctrlx]. We run all experiments on NVIDIA Tesla V100 GPUs.
@@ E.2 Dataset Details
We construct our dataset based on the conditional generation datasets from Ctrl-X [lin2024ctrlx] and FreeControl [mo2023freecontrol]. Specifically, for conditions canny edge, depth map, normal map, HED edge, and scribble drawing, we select condition-prompt pairs from both datasets and merge them. We collect a total of 15 condition images per condition and form 22 condition-prompt pairs for canny edge and 21 pairs for each of the remaining four conditions.
For human pose and segmentation map, since both original datasets contain limited examples, we supplement them by collecting additional human pose images from the web and segmentation masks from the ADE20K [Zhou_2017_CVPR] dataset. We obtain 15 images for each of these two conditions and pair them with text prompts using a combination of templates and hand annotation, resulting in 21 image-prompt pairs for human pose and 23 for segmentation mask.
@@ Appendix F Additional Experimental Results
ControlNet [zhang2023controlNet] 0.067 0.309 0.701 0.509 0.298 0.285
T2I-Adapter [mou2023t2i] 0.116 0.287 0.728 0.636 -0.050 0.261
SDEdit [meng2021sdedit] 0.154 0.259 0.315 0.734 -1.374 0.189
P2P [hertz2022prompt-to-prompt] 0.197 0.251 0.266 0.724 -1.786 0.168
PnP [tumanyan2023plug-and-play] 0.157 0.256 0.151 0.724 -1.789 0.168
InfEdit [xu2023infedit] 0.135 0.296 0.357 0.636 -0.202 0.244
FreeControl [mo2023freecontrol] 0.116 0.320 0.667 0.626 0.554 0.285
Ctrl-X [lin2024ctrlx] 0.104 0.315 0.650 0.579 0.291 0.283
@@ F.1 Inference Efficiency
Our framework inherits the computational efficiency of injection-based pipelines. Moreover, since $g(t)=C$ in our implementation, our pipeline is able to cache and reuse $\mathbf{f}_{l,g(t)}^{\text{struct}}$, in contrast to previous synchronous pipelines that perform inference at every timestep. Tab. 3 shows the average inference time of different settings in our experiments, with the refiner of SDXL [podell2023sdxl] enabled. As shown in the table, structure-rich injection saves inference time by up to 9% with no noticeable degradation in generation quality. Our framework still achieves superior inference efficiency even with the restart refinement schedule, demonstrating simultaneous improvements in both generation quality and inference speed. Appearance-rich prompting module improves the robustness of the framework at the expense of an 8% inference time overhead. Overall, this training-free framework provides users with the flexibility to adapt to task-specific priorities and negligible overhead in most cases.
@@ F.3 Additional Quantitative Results
Since T2I-Adapter-SDXL [mou2023t2i] supports only four (canny, depth, normal, and pose) out of the seven condition types in our full dataset, we further conduct a quantitative comparison limited to these four types. As shown in Tab. 2, our method outperforms all baselines across almost every metric. Notably, these metrics jointly assess both structure preservation (e.g., DINO self-similarity [tumanyan2022splicing], DreamSim [fu2023dreamsim]) and generation quality (e.g., ImageReward [xu2023imagereward], HPSv2 [wu2023hpsv2]), highlighting the effectiveness of our approach.